### v1

# Anti-bosonic current in thermal spin pumping

Abstract
We investigate thermal spin pumping in gadolinium iron garnet, focusing on dynamics of antiferromagnetic magnon and its impact on spin and heat transport. Antiferromagnet has right-handed and left-handed magnon modes. Using a two-magnetic-lattice model, we derive temperature-dependent spin-wave spectra and thermal spin pumping constants, revealing distinct transport regimes governed by competing contribution of magnon modes. We show that competition of both modes induces a compensation temperature in thermally generated spin current. This theoretical framework deepens the understanding of spin-caloritronic effects in ferrimagnets, offering new perspectives for spintronic device optimization.
Introduction. The exploration of fundamental particles has driven significant advancements in our understanding of quantum mechanics and condensed matter physics. Among these particles, the study of antiferromagnetic magnons, i.e., the quanta of spin waves in antiferromagnetic systems, plays a pivotal role in various theoretical frameworks, including quantum field theory and spintronics. Magnons are essential for understanding phenomena such as spin transport and the manipulation of spin states, which are crucial in quantum information science and spintronics Lebrun et al. (2020); Cornelissen et al. (2015). Recent advancements have demonstrated that antiferromagnets can be explored as promising systems for spintronic applications, including quantum computing and communication technologies Zhou et al. (2022); Poelchen et al. (2023).
Building on the concept of magnons, researchers have introduced the notion of magnons with different handedness, which are spin-wave excitations that can precess in opposite directions, depending on the underlying magnetic structure. In particular, the antiferromagnetic magnon in systems with alternating spin orientations exhibits different modes of propagation, including left-handed and right-handed modes Kanj et al. (2023), as illustrated in Fig. 1. Detecting antiferromagnetic magnons with distinct handedness represents an emerging frontier in condensed matter physics, with techniques such as thermal spin pumping offering promising potential Siddiqui et al. (2021). Thermal spin pumping utilizes thermally excited spin waves to generate electrical signals, providing an indirect method to probe magnonic properties. By integrating thermal spin pumping with advanced detection techniques, researchers could pave the way for groundbreaking discoveries in quantum magnonics and spin-based computing Ortiz et al. (2021). The coupling of thermal spin pumping with magnons could facilitate the manipulation and detection of spin waves, particularly those with unique handedness, thereby enhancing their applicability for future technological developments Holzmann et al. (2022).
Here, we proposed that the distinct chirality of the antiferromagnetic magnon modes can be conceptualized as bosonic quasiparticles analogous to how holes in semiconductor systems are considered antifermionic quasiparticles. This novel perspective allows us to view the handedness of magnons as a form of antiboson, thereby enriching the theoretical framework of magnonics Lachance-Quirion et al. (2019). The potential applications of these distinct modes also hold promise for innovative approaches in quantum magnonics and the manipulation of information via spintronic devices Weiss et al. (2023); Bai et al. (2015).
Figure 1: (a) Right-handed magnon precessing with radial frequency $\omega_{R}$ in counter clockwise direction around $+z$ axis and left-handed magnon precessing with radial frequency $\omega_{L}$ in clockwise direction.
Rare earth iron garnets (RIGs) are particularly well-suited for studying antiferromagnetic magnons due to their unique magnetic properties and exceptionally low magnetic damping A. Seo et al. (2024). The intricate magnetic structure of RIGs—including three magnetic sublattices (tetrahedral Fe3+ sites, octahedral Fe3+ sites, and dodecahedral rare earth R3+ sites)—generates a rich spin wave spectrum with multiple branches Barker and Bauer (2016); Huang et al. (2024). Tuning the rare earth ion composition allows for precise engineering of the spin wave spectrum, potentially facilitating the detection of magnons with distinct handedness Zhang et al. (2023). Furthermore, RIGs, particularly yttrium iron garnet (YIG), exhibit exceptionally low magnetic damping, rendering them ideal for studying spin wave dynamics and the integration of spintronic devices Cornelissen et al. (2015); Li et al. (2020). The presence of rare earth ions in RIGs introduces strong spin-orbit coupling effects, which may induce phenomena such as the inverse Faraday effect and the Dzyaloshinskii-Moriya interaction, critical for the manipulation and detection of spin waves with distinct handedness Cheng et al. (2016).
YIG thermal spin pumping. In thermal spin pumping, spin $J_{s}$ and heat $J_{Q}$ currents are transferred from a magnetic material, such as rare-earth iron garnet, to a non-magnetic metal, such as Pt. This spin current is then converted into an electrical signal via the inverse spin Hall effect. At the interface, $J_{s}$ and $J_{Q}$ coupled and can be expressed through a linear matrix relation involving the driving forces: spin accumulation $\bm{\mu}_{s}$ and temperature difference $\Delta T$, analogous to the thermoelectric effect:
$\left[\begin{array}[]{c}J_{s}\\ J_{Q}\end{array}\right]=\left[\begin{array}[]{cc}\mathcal{L}_{0}&\mathcal{L}_{1}\\ \mathcal{L}_{1}&\mathcal{L}_{2}\end{array}\right]\left[\begin{array}[]{c}-\hat{\textbf{m}}\cdot\bm{\mu}_{s}\\ -\Delta T/T\end{array}\right].$ (1)
Here, $\hat{\textbf{m}}=z\hat{z}S_{z}/|S_{z}|$ is the average direction of the magnetization. $\mathcal{L}_{i}$ is analogous to thermoelectric transport coefficients. $\mathcal{L}_{0}$ represents the interface spin injection conductance, while $\mathcal{L}_{2}$ denotes the heat conductance. The ratios $\mathcal{L}_{1}/\mathcal{L}_{0}$ and $\mathcal{L}_{1}/(\mathcal{L}_{0}T)$ correspond to the spin Peltier and spin Seebeck coefficients, respectively Cahaya et al. (2014).
The expression of $\mathcal{L}_{0}$ and $\mathcal{L}_{1}$ can be determined from total spin current transferred at the interface.
$J_{s}=$ $J_{\mathrm{bf}}+J_{\mathrm{sp}}+J_{\mathrm{fl}},$ (2)
where $J_{\mathrm{sp}}$ arises from the spin pumping
$J_{\mathrm{sp}}=$ $G\left<\textbf{S}\times\dot{\textbf{S}}\right>\cdot\hat{z}.$ (3)
$J_{\mathrm{fl}}=-\gamma\left<\textbf{S}\times\textbf{h}^{\prime}\right>\cdot\hat{z}$ is spin backflow due to thermal noise. $J_{\mathrm{fl}}=-J_{\mathrm{sp}}(T-\Delta T)/T$ has been shown to balances $J_{\mathrm{sp}}$ when there is no temperature difference across interface Xiao et al. (2010).
$J_{\mathrm{sp}}+J_{\mathrm{fl}}=$ $\frac{\Delta T}{T}G\left<\textbf{S}\times\dot{\textbf{S}}\right>\cdot\hat{z}.$ (4)
Lastly, $J_{\mathrm{bf}}$ arises from spin backflow due to spin accumulation polarized along the $z$-direction:
[…]
where $G$ is the spin-mixing conductanceJiao and Bauer (2013).
Using the Holstein-Primakoff transformation $S_{+}=S_{x}+iS_{y}\simeq\sqrt{2S}a$, $S_{-}=S_{x}-iS_{y}\simeq\sqrt{2S}a^{\dagger}$, $S_{z}=S-a^{\dagger}a$, we obtain
[…]
The expectation value of the bosonic field of the ferromagnetic spin wave $\omega=\omega_{0}+Dk^{2}$ gives
$\mathcal{L}_{0}=$ $2GS_{z}\int_{\omega_{0}}^{\infty}\frac{\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}=\frac{GS_{z}}{2}\left(\frac{k_{B}T}{\pi D}\right)^{\frac{3}{2}}\mathrm{Li}_{\frac{3}{2}}\left(e^{-\beta\omega_{0}}\right),$ (7)
where $\beta=(k_{B}T)^{-1}$, $\mathrm{Li}_{n}(x)$ is the polylogarithm function, with $\mathrm{Li}_{n}(1)=\zeta(n)$ reducing to the Riemann zeta function. The rest of the spin current in Eq. 2 is contributes to $\mathcal{L}_{1}$
[…]
$=2GS_{z}\int_{0}^{\infty}\frac{\omega\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}=\frac{3GS_{z}\left(k_{B}T\right)^{\frac{5}{2}}}{4\left(\pi D\right)^{\frac{3}{2}}}\mathrm{Li}_{\frac{5}{2}}\left(e^{-\beta\omega_{0}}\right).$ (8)
In YIG, the magnon gap $\omega_{0}=\gamma H$ can be non zero due to Zeeman effect Rodrigue et al. (1960). Since $\lim_{x\to 0}\mathrm{Li}_{n}(x)\simeq x$, the magnetic field $H$ exponentially suppress thermal spin pumping Kikkawa et al. (2015) $\mathcal{L}_{1}\propto e^{-\gamma H/k_{B}T}.$ The magnetization of YIG originates from five Fe3+ ions distributed across two magnetic lattices. Two of these ions occupy tetrahedral Fe3+ sites, while three are in octahedral sites. Due to their antiferromagnetic coupling, only one Fe3+ spin contributes Néel et al. (1964), as illustrated in Fig. 2a. When yttrium in YIG is substituted by lanthanides, R3+ occupies the dodecahedral site, forming an additional magnetic lattice. This substitution leads to a more complex spin-wave spectrum compared to that of YIG. While significant computational efforts are underway to unravel these spectra Wang et al. (2020); Li et al. (2022), determining thermal spin pumping from detailed spectra remains a formidable challenge Shen (2019). Additionally, conduction electrons interact differently with localized spin and orbital angular momentum Kondo (1962). To focus on competition of antiferromagnetic magnon due to ferrimagnetically coupled spins, we restrict our discussion to gadolinium iron garnet (GdIG), which exhibits zero orbital angular momentum.
Statistical mechanics of GdIG. A model that provides considerable insight into thermal spin pumping in GdIG is the two-magnetic-lattice model Cahaya (2022); Cahaya et al. (2022). In this model, the magnon spectrum of GdIG is determined by two magnetic lattices: $S^{\mathrm{Gd}}$, representing the total angular momentum of three Gd3+ ions, and $S^{\mathrm{Fe}}$, representing the spin angular momentum of Fe3+ ions. This model has been shown to describe both its low-lying spin wave spectrum Tinkham (1961) and room-temperature thermal spin pumping Cahaya (2022).
We consider two isotropic exchange couplings: Fe3+-Fe3+ coupling, characterized by the exchange constant $I$, and Fe3+-Gd3+ coupling, characterized by the exchange constant $\lambda$, as illustrated in Fig. 2b.
$\mathcal{H}=$ $-\frac{I}{2}\sum_{n}\textbf{S}^{\mathrm{Fe}}_{j}\cdot\textbf{S}^{\mathrm{Fe}}_{j+1}+\lambda\sum_{j}\textbf{S}^{\mathrm{Gd}}_{j}\cdot\textbf{S}^{\mathrm{Fe}}_{n},$ (9)
where the subscript $n$ indexes the Fe3+ chain. The Fe-Gd coupling induces a paramagnetic response of $S^{\mathrm{Gd}}$ to the molecular field generated by $S^{\mathrm{Fe}}$. Gd-Gd coupling is neglected, as validated by the dominance of Fe-Fe coupling in the Curie temperature $T_{C}=IS^{\mathrm{Fe}}(S^{\mathrm{Fe}}+1)/(3k_{B})+\mathcal{O}\left(\lambda^{2}/I\right)$, Gilleo (1980) obtained from the statistical average:
$S^{\mathrm{Gd}}_{z}=$ $3S^{\mathrm{Gd}}B_{S^{\mathrm{Gd}}}\left(S^{\mathrm{Gd}}\frac{-\lambda S^{\mathrm{Fe}}_{z}}{k_{B}T}\right)\simeq-\frac{S^{\mathrm{Gd}}(S^{\mathrm{Gd}}+1)\lambda S^{\mathrm{Fe}}_{z}}{k_{B}T},$
$S^{\mathrm{Fe}}_{z}=$ $S^{\mathrm{Fe}}B_{S^{\mathrm{Fe}}}\left(S^{\mathrm{Fe}}\frac{IS^{\mathrm{Fe}}_{z}-\lambda S^{\mathrm{Gd}}_{z}}{k_{B}T}\right),$ (10)
where $B_{J}(x)$ is the Brillouin function. Fig. 2c illustrates the temperature dependence of the spins. A compensation temperature occurs at Cahaya (2022)
$T_{B}=k_{B}^{-1}\lambda S^{\mathrm{Gd}}(S^{\mathrm{Gd}}+1).$ (11)
Figure 2: (a) Exchange coupling between spins of Fe3+, occupying tetrahedral and octahedral sites, and Gd3+, occupying dodecahedral sites, in Gd3Fe5555O1212{12}12 (GdIG) can be approximated by (b) the two-spin model with $\mathrm{S}^{\mathrm{Fe}}$ and $\mathrm{S}^{\mathrm{Gd}}$. (c) Temperature dependence of the $z$-component of spins in GdIG. The total spin is zero at the compensation temperature $T_{B}$.
Spin-wave spectrum. The interaction leads to the following Landau–Lifshitz equations of motion for the spins
$\frac{d\textbf{S}^{\mathrm{Fe}}_{j}}{dt}=$ $\textbf{S}^{\mathrm{Fe}}_{n}\times\left(\frac{I}{2}\left(\textbf{S}^{\mathrm{Fe}}_{j+1}+\textbf{S}^{\mathrm{Fe}}_{n-1}\right)-\lambda\textbf{S}^{\mathrm{Gd}}_{j}\right)$
$\frac{d\textbf{S}^{\mathrm{Gd}}_{j}}{dt}=$ $-\lambda\textbf{S}^{\mathrm{Gd}}_{j}\times\textbf{S}^{\mathrm{Fe}}_{j},$ (12)
The spin-wave spectrum is evaluated by linearizing the Landau-Lifshitz equations using $F_{jx}\pm F_{jy}=F_{\pm}e^{i\left(jka-\omega t\right)}$
$i\frac{d}{dt}\left[\begin{array}[]{c}S^{\mathrm{Fe}}_{+}\\ S^{\mathrm{Gd}}_{+}\end{array}\right]=W\left[\begin{array}[]{c}S^{\mathrm{Fe}}_{+}\\ S^{\mathrm{Gd}}_{+}\end{array}\right].$ (17)
$W=\left[\begin{array}[]{cc}IS^{\mathrm{Fe}}_{z}(1-\cos ka)-\lambda S^{\mathrm{Gd}}_{z}&\lambda S^{\mathrm{Fe}}_{z}\\ \lambda S^{\mathrm{Gd}}_{z}&-\lambda S^{\mathrm{Fe}}_{z}\end{array}\right].$ (20)
For small $k$, the eigen frequencies are
$\omega_{1}=$ $\frac{S^{\mathrm{Fe}}_{z}D_{\mathrm{YIG}}k^{2}}{S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}},$
$\omega_{2}=$ $-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)+\frac{S^{\mathrm{Gd}}_{z}D_{\mathrm{YIG}}k^{2}}{S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}},$ (21)
where $D_{0}=Ia^{2}S^{\mathrm{Fe}}_{z}/2$ is the spin wave stiffness of YIG. The eigenfrequencies $\omega_{R}=\mathrm{max}\left(\omega_{1},\omega_{2}\right)$ and $\omega_{L}=\mathrm{min}\left(\omega_{1},\omega_{2}\right)$ are associated with right-handed (R) and left-handed (L) magnons, respectively. Their precession directions are illustrated in Fig. 1. Since $S^{\mathrm{Gd}}_{z}$ varies according to Eq. 10, the spectrum changes with temperatures, as illustrated in Fig. 3.
Figure 3: Spin-wave spectrum of GdIG at temperatures (a) lower than, (b) equal to, and (c) higher than the compensation temperature $T_{B}$. Positive frequency ($\omega_{R}$) is associated with right-handed magnon. On the other hand, negative frequency ($\omega_{L}$), with a minimum frequency $\omega_{\mathrm{min}}\simeq-\lambda S^{\mathrm{Fe}}_{z}$, is associated with right-handed magnonboth magnon branches.
$\mathcal{L}_{0}$ and $\mathcal{L}_{1}$ can be modified to account for the both magnon branches by substituting $S_{z}=S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}$ and
$\int_{-\infty}^{\infty}\frac{\omega^{n}\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}$
$=\int_{0}^{\infty}\frac{\omega_{R}^{n}\mathrm{DOS}(\omega_{R})d\omega_{R}}{e^{\beta\omega_{R}}-1}+\int_{-\infty}^{0}\frac{\omega_{L}^{n}\mathrm{DOS}(|\omega_{L}|)d|\omega_{L}|}{e^{\beta|\omega_{L}|}-1}$
For negative energy spectra, one can see that the Bose-Einstein distribution become
$\frac{1}{e^{\beta|\omega_{L}|}-1}=\frac{1}{e^{-\beta\omega_{L}}-1}.$ (22)
Thus, the contributions of R and L-magnons are additive in $\mathcal{L}_{0}$
$\mathcal{L}_{0}=$ $2G\left|S_{z}^{\mathrm{Fe}}+S_{z}^{\mathrm{Gd}}\right|\left(\int_{0}^{\infty}\frac{\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}+\int_{-\infty}^{0}\frac{\mathrm{DOS}(\omega)d\omega}{e^{-\beta\omega}-1}\right)$
$\mathcal{L}_{1}=$ $2G\left(S_{z}^{\mathrm{Fe}}+S_{z}^{\mathrm{Gd}}\right)\left(\int_{0}^{\infty}\frac{\omega\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}+\int_{-\infty}^{0}\frac{\omega\mathrm{DOS}(\omega)d\omega}{e^{-\beta\omega}-1}\right)$
but subtractive in $\mathcal{L}_{1}$, analogous to the transport of electrons and holes in semiconductors (see Table 1). While band structure of semiconductor is weakly depends on temperature for as long as $k_{B}T\ll\epsilon_{F}$ is a lot smaller than the Fermi energy $\epsilon_{F}$, the magnon spectrum of ferrimagnetic Gd3Fe5O12 is strongly depends on temperature, as illustrated in Fig. 3)
Table 1: Comparison of thermal spin pumping from ferrimagnetic magnons to thermoelectric transport in semiconductor.
Gd3Fe5O12 Semiconductor
Boson Fermion
$R$-magnon ($\omega_{R}$ branch) Electron (conduction band)
$n_{R}\propto\displaystyle\int\frac{\mathrm{DOS}(\omega_{R})d\omega_{R}}{e^{\beta\omega_{R}}-1}$ $n_{e}\propto\displaystyle\int\frac{\mathrm{DOS}(\epsilon)d\epsilon}{e^{\beta(\epsilon-\epsilon_{F})}+1}$
$L$-magnon ($\omega_{L}$ branch) Hole (valence band)
$n_{L}\propto\displaystyle\int\frac{\mathrm{DOS}(\omega_{L})d\omega_{L}}{e^{-\beta\omega_{L}}-1}$ $n_{h}\propto\displaystyle\int\frac{\mathrm{DOS}(\epsilon)d\epsilon}{e^{-\beta(\epsilon-\epsilon_{F})}+1}$
$\mathcal{L}_{0R}\mathcal{L}_{0L}>0$ $\mathcal{L}_{0e}\mathcal{L}_{0h}>0$
$\mathcal{L}_{1R}\mathcal{L}_{1L}<0$ $\mathcal{L}_{1e}\mathcal{L}_{1h}<0$
High temperature spin pumping. We first consider temperatures higher than the compensation temperature $T>T_{B}$. In this regime, the net magnetization $S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}$ is positive. Thus $\omega_{R}=\omega_{1}$ and $\omega_{L}=\omega_{2}$. R-magnon dominate spin conductance at high temperatures because the number of L-magnon remains significantly smaller.
$\left(\frac{n_{L}}{n_{R}}\right)_{T>T_{B}}=\left(\frac{\mathcal{L}_{0L}}{\mathcal{L}_{0R}}\right)_{T>T_{B}}=\frac{\int_{\omega_{\mathrm{min}}}^{-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)}\frac{\mathrm{DOS}\left(\omega_{L}\right)d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{0}^{\infty}\frac{\mathrm{DOS}\left(\omega_{R}\right)d\omega_{R}}{e^{\beta\omega_{R}}-1}}$
$\simeq\frac{4\left|\frac{S_{z}^{\mathrm{Fe}}}{S_{z}^{\mathrm{Gd}}}\right|^{\frac{3}{2}}}{\zeta\left(\frac{3}{2}\right)}\int^{-\beta\lambda S^{\mathrm{Gd}}_{z}}_{0}\frac{(2\pi^{2})^{-1}x^{\frac{1}{2}}dx}{e^{\beta\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)+x}-1}\ll 1.$ (23)
Here $\omega_{\mathrm{min}}\simeq-\lambda S^{\mathrm{Fe}}_{z}$ is the minimum frequency of the L-magnon spectrum, as illustrated in Fig. 3. Thus, the R-magnon contribution dominates in high-temperature spin conductance.
In this temperature regime, the L-magnon contribution is still much smaller
$\left(\frac{\mathcal{L}_{1L}}{\mathcal{L}_{1R}}\right)_{T>T_{B}}=\frac{\int_{\omega_{\mathrm{min}}}^{-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)}\frac{\omega_{L}\mathrm{DOS}\left(\omega_{L}\right)d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{0}^{\infty}\frac{\omega_{R}\mathrm{DOS}\left(\omega_{R}\right)d\omega_{R}}{e^{\beta\omega_{R}}-1}}$
$\simeq\frac{-8\left|\frac{S_{z}^{\mathrm{Fe}}}{S_{z}^{\mathrm{Gd}}}\right|^{\frac{3}{2}}}{3\zeta\left(\frac{5}{2}\right)}\int^{-\beta\lambda S^{\mathrm{Gd}}_{z}}_{0}\frac{\left[\beta\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)+x\right]x^{\frac{1}{2}}dx}{2\pi^{2}\left[e^{\beta\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)+x}-1\right]}\ll 1.$ (24)
Thus, R-magnon dominate $\mathcal{L}_{1}$ and the spin Seebeck coefficient $\mathcal{L}_{1}/(\mathcal{L}_{0}T)$ at high temperatures, as depicted in Fig. 4. We can see that R and L-magnons contributions are substractive in $\mathcal{L}_{1}$.
Figure 4: (a) The interface spin injection conductance $\mathcal{L}_{0}$.R-magnon dominates transport at high temperatures, while L-magnon becomes dominant at low temperatures. (b) The thermal spin current $\mathcal{L}_{1}$ exhibits an additional compensation temperature $T_{A}$, where L-magnon contribution exactly cancels R-magnon contribution. (c) The ratio $\mathcal{L}_{1}/(\mathcal{L}_{0}T)$ determine the measured spin Seebeck voltage ($V_{\mathrm{SSE}}$). Our theoretical result is in agreement with experiment data from Ref. Cramer et al., 2017.
Low temperature spin pumping. We now consider temperatures lower than the compensation temperature $T<T_{B}$, where $\omega_{R}=\omega_{2}$ and $\omega_{L}=\omega_{1}$. Thus, the magnon gap shifts to $\omega_{R}$. Thus, reducing number of R-magnon
$\left(\frac{n_{L}}{n_{R}}\right)_{T<T_{B}}=\left(\frac{\mathcal{L}_{0L}}{\mathcal{L}_{0R}}\right)_{T<T_{B}}=\frac{\int_{\omega_{\mathrm{min}}}^{0}\frac{\mathrm{DOS}(\omega_{L})d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)}^{\infty}\frac{\mathrm{DOS}(\omega_{R})d\omega_{R}}{e^{\beta\omega_{R}}-1}}$
$=\frac{\frac{2}{\sqrt{\pi}}\left|\frac{S_{z}^{\mathrm{Gd}}}{S_{z}^{\mathrm{Fe}}}\right|^{\frac{3}{2}}}{\mathrm{Li}_{\frac{3}{2}}\left(e^{\beta\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)}\right)}\int^{\beta\lambda S^{\mathrm{Fe}}_{z}}_{0}\frac{x^{\frac{1}{2}}dx}{e^{x}-1}$
As the temperature decreases, the L-magnon population increases, making it the dominant contribution at low temperatures. Figure 4a shows the variation of $\mathcal{L}_{0}$ with temperature. At low temperatures, L-magnon dominate the spin conductance, with the asymptotic low-temperature limit
$\lim_{T\to 0}\mathcal{L}_{0}=\frac{G\zeta\left(\frac{3}{2}\right)\left(k_{B}T\right)^{{3}/{2}}\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|^{{5}/{2}}}{2\left(\pi D_{0}S^{\mathrm{Fe}}_{z}\right)^{{3}/{2}}}.$ (25)
Similarly, the R-magnon contribution to $\mathcal{L}_{1}$ is exponentially suppressed
$\left(\frac{\mathcal{L}_{1L}}{\mathcal{L}_{1R}}\right)_{T<T_{B}}=\frac{\int_{\omega_{\mathrm{min}}}^{0}\frac{\omega_{L}\mathrm{DOS}(\omega_{L})d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{-\lambda\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|}^{\infty}\frac{\omega_{R}\mathrm{DOS}(\omega_{R})d\omega_{R}}{e^{\beta\omega_{R}}-1}}$
$\simeq\frac{-\frac{4}{\sqrt{\pi}}\left|\frac{S_{z}^{\mathrm{Gd}}}{S_{z}^{\mathrm{Fe}}}\right|^{\frac{3}{2}}}{3\mathrm{Li}_{\frac{5}{2}}\left(e^{-x_{0}}\right)-2x_{0}\mathrm{Li}_{\frac{3}{2}}\left(e^{-x_{0}}\right)}\int^{\beta\lambda S^{\mathrm{Fe}}_{z}}_{0}\frac{x^{\frac{3}{2}}dx}{e^{x}-1}.$ (26)
Here $x_{0}=-\beta\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)$. A compensation temperature $T_{A}$ occurs when the total $\mathcal{L}_{1R}$ and $\mathcal{L}_{1L}$ cancel each others, as illustrated in Fig.4b. The value of $\mathcal{L}_{1}$ becomes negative for $T_{A}<T<T_{B}$.Numerical evaluation of the integrals yields $T_{A}/T_{B}=0.3$, which aligns well with experimental values reported in Ref. Geprägs et al., 2016 ($T_{A}=80$ and $T_{B}=256.5$) and Ref. Cramer et al., 2017 ($T_{A}=72$ and $T_{B}\simeq 250$).
Conclusion. In summary, we explored the mechanisms behind thermal spin pumping in GdIG, emphasizing the competition of right-handed and left-handed magnons. As magnon gap appears at low temperature, R-magnon transport is suppressed and replaced by L-magnon. The theoretical framework developed highlights the crucial roles of antibosonic spin-wave dispersion in determining thermal spin pumping.
By deriving the temperature-dependent spin wave spectrum of GdIG and its influence on spin and heat response constants ($\mathcal{L}_{0}$ and $\mathcal{L}_{1}$), we demonstrated how antiferromagnetic magnons dynamics influence interface spin conductance and spin Seebeck coefficient. Our analysis underscores the significance of temperature-dependent magnons contributions, highlighting their competitive effects in thermal spin pumping. We identified distinct transport regimes where R-magnon dominate at high temperatures, while L-magnon become relevant at low temperatures. This dual contribution leads to notable features such as compensation points where thermal spin currents vanish. The compensation temperatures thermal spin pumping of GdIG were quantitatively captured.
The developed theoretical approach offers a comprehensive framework to understand thermal spin pumping phenomena in complex magnetic materials. These results can serve as a foundation for future experimental and computational studies aimed at optimizing spintronic devices based on thermal spin transport.
Acknowledgments. We acknowledge the support of Universitas Indonesia via PUTI Research Grant No. NKB-366/UN2.RST/HKP.05.00/2024